For each of the following, compute the present value (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.):

Present Value Years   Interest Rate Future value
$         14       8 %   $ 16,151  
        5       14       58,557  
        30       15       893,073  
        35       8       557,164  
Answer #1

To calculate Present value we use the following formulae

PV=C/(1+i)n

where PV is present value,C is future value ,I is rate of interest and n is no of years

Case 1)PV=16,151/(1+0.08)14=16,151/(1.08)14=16,151/2.937=$5499.15

case2)PV=58,557/(1+0.14)5=58,557/(1.14)5=58,557/1.925=$30,419.22

case3) PV=893,073/(1+0.15)30=893,073/(1.15)30=893,073/66.211=$13,488.29

case4)PV=557,164(1+0.08)35=557,164/(1.08)35=557,164/14.785=$37,684.41